OpenStudy (andijo76):

Statistics question Please help I am not sure if i am right or on the right track...

OpenStudy (andijo76):

Express the confidence interval (0.042,0.134) in the form of p^-E<p^+E phat= (0.042+0.134)/2 phat=(0.176)2 phat=0.088 E=0.134-0,088 E=0.046 0.088<p<0.046

OpenStudy (andijo76):

correct?

OpenStudy (tkhunny):

No. You went past your mark. Find p-hat (0.042+0.134)/2 = 0.176/2 = 0.088 -- Perfect. Find E (0.134-0.042)/2 = 0.092/2 = 0.046 0.088 - 0.046 < 0.088 + 0.046 Checking, just for fun. 0.088 - 0.046 = 0.042 Good 0.088 + 0.046 = 0.134 Good